# Adaptive Informed Deep Neural Networks for Power Flow Analysis

### v2

@@ V-B5 Scalability
The experiments are extended to compare the performance of PINN4PF, MLP, and LR across different test system sizes: 4-bus, 15-bus, 290-bus, and 2224-bus test systems. The maximum testing MSE for the direct physical quantities obtained for all test system sizes is presented in Table II. Accordingly, PINN4PF significantly outperforms MLP and LR. Fig. 6 compares the performance of PINN4PF and MLP in terms of direct physical quantities under extreme conditions for the 2224-bus test system. PINN4PF achieves a maximum MSE $[V^{2}]$ 50% lower than MLP for this test system. This improvement is crucial for power system operations, particularly during unexpected events that cause deviations in $v$ from the nominal value. The analysis also considers the performance of PINN4PF, MLP, and LR in approximating derived physical quantities. The maximum testing MSE for all test system sizes is presented in Table III. The table demonstrates the superiority of PINN4PF over MLP and LR across all test system sizes. For the 2224-bus test system, for example, the maximum MSE $[A^{2}]$ obtained by PINN4PF is respectively 92% and 98% lower than MLP and LR. In addition, as the size of the test system increases, the performance gap between PINN4PF, MLP, and LR becomes more pronounced. Fig. 7 compares the performance of PINN4PF and MLP under extreme conditions in the 2224-bus test system for the derived physical quantity. It should be noted that while the testing MSE for the direct quantities, i.e., voltage magnitude and phase angle, shows marginal differences between PINN4PF and MLP, PINN4PF significantly outperforms MLP in terms of the derived physical quantities, such as line current.
@@
TABLE II: Performance comparison of power flow solvers based on MSE obtained for direct physical quantities.
Case Model MSE [$V^{2}$] MSE [$rad^{2}$]
@@
TABLE III: Performance comparison of power flow solvers based on MSE obtained for derived physical quantities.
Case Model MSE [$A^{2}$] MSE [$W^{2}$] MSE [$VAR^{2}$]
@@
2224-bus PINN4PF $1.15\times 10^{-6}$ $1.13\times 10^{-8}$ $9.45\times 10^{-7}$
@@
V-C Model Sensitivity Analysis
@@ V-C1 Sensitivity analysis for MLP
V-C2 Sensitivity analysis for PINN4PF
@@
V-C3 Impact of modifications on the performance of PINN4P
@@
TABLE IV: Performance comparison of different configurations of PINN4PF based on MSE obtained for direct physical quantities and the 15-bus test system.
Model MSE [$V^{2}$] MSE [$rad^{2}$]
@@
VI Discussion
The results of this study demonstrate the effectiveness and robustness of the proposed PINN4PF architecture in power flow analysis. Several key points and observations arise from the findings:
@@ VI Discussion
VII Conclusion

### v3

@@ 5.2.5 Scalability
The experiments are extended to compare the performance of PINN4PF, MLP, and LR across different test system sizes: 4-bus, 15-bus, 290-bus, and 2224-bus test systems. The maximum testing MSE for the direct physical quantities obtained for all test system sizes is presented in Table II. Accordingly, PINN4PF significantly outperforms MLP and LR. Fig. 6 compares the performance of PINN4PF and MLP in terms of direct physical quantities under extreme conditions for the 2224-bus test system. PINN4PF achieves a maximum MSE $V^{2}]$ 50% lower than MLP for this test system. This improvement is crucial for power system operations, particularly during unexpected events that cause deviations in $v$ from the nominal value. The analysis also considers the performance of PINN4PF, MLP, and LR in approximating derived physical quantities. The maximum testing MSE for all test system sizes is presented in Table III. The table demonstrates the superiority of PINN4PF over MLP and LR across all test system sizes. For the 2224-bus test system, for example, the maximum MSE $A^{2}]$ obtained by PINN4PF is respectively 92% and 98% lower than MLP and LR. In addition, as the size of the test system increases, the performance gap between PINN4PF, MLP, and LR becomes more pronounced. Fig. 7 compares the performance of PINN4PF and MLP under extreme conditions in the 2224-bus test system for the derived physical quantity. It should be noted that while the testing MSE for the direct quantities, i.e., voltage magnitude and phase angle, shows marginal differences between PINN4PF and MLP, PINN4PF significantly outperforms MLP in terms of the derived physical quantities, such as line current.
@@
Table 2: Performance comparison of power flow solvers based on MSE obtained for direct physical quantities.
Case Model MSE $V^{2}$] MSE $rad^{2}$]
@@
Table 3: Performance comparison of power flow solvers based on MSE obtained for derived physical quantities.
Case Model MSE $A^{2}$] MSE $W^{2}$] MSE $VAR^{2}$]
@@
2224-bus PINN4PF $15\times 10^{-6} 1.1$$13\times 10^{-8}$ $9.45\times 10^{-7}$
@@
5.3 Model Sensitivity Analysis
@@ 5.3.1 Sensitivity analysis for MLP
5.3.2 Sensitivity analysis for PINN4PF
@@
5.3.3 Impact of modifications on the performance of PINN4P
@@
Table 4: Performance comparison of different configurations of PINN4PF based on MSE obtained for direct physical quantities and the 15-bus test system.
Model MSE $V^{2}$] MSE $rad^{2}$]
@@
6 Discussion
Several key points and observations arise from the findings:
  • 1.
@@ 6 Discussion
  • 6.
    While PINN4PF demonstrates strong performance under fixed network topology, it requires retraining if the system topology changes, as the current architecture does not support topological generalization. Future research may explore the potential of transfer learning to adapt pre-trained models to new topologies with minimal additional training. Note, however, that although recent advances in GNNs aim to support topological changes, they still do not provide consistently reliable results at the time of writing this paper. Therefore, enabling robust topological adaptability remains an open research challenge for NNs.
7 Conclusion
@@
Acknowledgment
This work is part of the DATALESs project (project no. 482.20.602), which is jointly financed by Netherlands Organization for Scientific Research (NWO), and National Natural Science Foundation of China (NSFC). The work used the Dutch national e-infrastructure with the support of the SURF Cooperative (grant no. EINF-6569).